OpenStudy (joshoyen):

Factor this equation. 1. 4x^2 + 7x + 3 = 0

OpenStudy (sloppycanada):

x = -1 or x = 3/4

OpenStudy (dayakar):

Split the middle term

OpenStudy (dayakar):

4x^2+4x+3x+3

OpenStudy (joshoyen):

thanks guys

OpenStudy (dayakar):

welcome

OpenStudy (anonymous):

\[4 x^2+7 x+3=(x+1) (4 x+3)\]